OpenStudy (anonymous):

Every new customer at the Burger Barn qualifies to receive a free whistle. The bag of whistles has 9 black ones, 6 red ones, 7 green ones, 10 pink ones, and 4 purple ones.

OpenStudy (anonymous):

and the question?

OpenStudy (anonymous):

What is the probability the next customer will receive a red one @answereverything

OpenStudy (anonymous):

the number of reds /the sum

OpenStudy (anonymous):

what do you mean? @answereverything

OpenStudy (anonymous):

use the number of reds divided by the total number of whistles

OpenStudy (anonymous):

@answereverything sorry im confused

OpenStudy (anonymous):

Probability Of An Event \[P(A)={ The Number Of Ways Event A Can Occur \over The Total Number Of Possible Outcomes}\]

OpenStudy (anonymous):

ok, first of all, there are 9+6+7+10+4=36 whisltes, but there is only 6 red whistles, so the chance of getting of red one is the number of reds/the number of total whistles

OpenStudy (anonymous):

In this case, the number of ways the event can occur is 6 (because it is the number of red whistles.) The total number of possible outcomes is the total of all the whistles in the bag. (Add everything up, including the red whistles.)

OpenStudy (anonymous):

\[6 \over 9+6+7+10+4\]

OpenStudy (anonymous):

then simplify.

OpenStudy (anonymous):

0.0625

OpenStudy (anonymous):

Leave it in fraction form.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Probability of Red = Number of red colored whistles / Total number of whistles = 6 / 36 = 1/6

OpenStudy (anonymous):

so its 1/6 the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!:)

OpenStudy (anonymous):

you're welcome.